# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 1.3 Characterization of Traces
$R_{n-1}(X+n-\tfrac{1}{2})-R_{n-1}(X-\tfrac{1}{2})=\sum_{i=0}^{n-1}P_{i}(X+i).$
Then $R_{0}$, …, $R_{n-2}$ are uniquely defined by the first $n-1$ components of the equation, and the resulting solution is valid because the sum $\sum_{i=0}^{n-1}(R_{i}(X+i+\frac{1}{2})-R_{i-1}(X+i-\frac{1}{2}))$ is equal to $R_{n-1}(X+n-\frac{1}{2})-R_{n-1}(X-\frac{1}{2})$, which equals $\sum_{i=0}^{n-1}P_{i}(X+i)$, as needed. We also note that the kernel of $\varphi$ is $\{(z,\dots,z):z\in\mathbb{C}\}$.
Letting
$U=\{(R_{0}(X),\ \dots,\ R_{n-1}(X)):R_{i}(X)=R_{-i-1}(-X)\},$
we claim that the preimage of $W$ under $\varphi$ is $U$. Indeed, $\varphi(U)\subseteq W$ because for any element $(R_{0}(X),\dots,R_{n-1}(X))\in U$, the $i$th component of $\varphi(R_{0}(X),\dots,R_{n-1}(X))$ is
$R_{i}(X+\tfrac{1}{2})-R_{i-1}(X-\tfrac{1}{2})=R_{-i-1}(-X-\tfrac{1}{2})-R_{-i}(-X+\tfrac{1}{2}),$
which is equal to the negative of the $(n-i)$th component. Conversely, for any element $(R_{0}(X),\dots,R_{n-1}(X))$ mapped to $W$ by $\varphi$, the element of $V$ with $i$th component equal to $\frac{R_{i}(X)+R_{-i-1}(-X)}{2}$ belongs to $U$ and is also mapped to $W$ by $\varphi$. Therefore,
$V/\langle\varphi(I),W\rangle\cong V/\langle I,U\rangle.$
Let $\psi:V\to V/I\cong\mathbb{C}^{n}$ be the quotient map given by mapping $(R_{0}(X),\dots,R_{n-1}(X))$ to $(R_{0}(\alpha_{0}),\dots,R_{n-1}(\alpha_{n-1}))$. Because $\alpha_{i}=-\alpha_{-i-1}$, $\psi(U)$ is the set $S$ of vectors $(v_{0},\dots,v_{n-1})$ such that $v_{i}=v_{-i-1}$ for all $i$. Therefore,
$V/\langle I,U\rangle\cong\mathbb{C}^{n}/S,$
the dimension of which is $\frac{n}{2}$, as desired. ∎
1.4 Analytic Formula for Traces
We now give an analytic description of traces using the notion of a good contour from Kly (23). In particular, similar to Definition 3.6 of Kly (23), we say that a non-self-intersecting curve $C$ with exactly one point in the set $\mathbb{R}+ri$ for all real $r$ is a good contour if the following three conditions hold:
  1. 1.
    There exist real $r>0$ and $a$ such that, letting $B_{r}$ be the disk of radius $r$ centered at the origin, $C\setminus B_{r}$ coincides with $(a+i\mathbb{R}_{>0})\cup(-a+i\mathbb{R}_{<0})\setminus B_{r}$. This property ensures that the notions of “to the left of $C$” and “to the right of C” are well-defined.
  2. 2.
@@ 1.4 Analytic Formula for Traces
  3. 3.
[…]
Because the parameter $c$ is assumed to be generic, no two $\alpha_{q}$ have the same imaginary part, so a curve satisfying Condition 222 exists. Also, since $\alpha_{q}=-\alpha_{-q-1}$ by Lemma 1.2.3, the curve may be taken to also satisfy Condition 333. Finally, since there are finitely many $\alpha_{q}$, the curve may also be taken to satisfy Condition 111, so a good contour is guaranteed to exist.
@@
$\textbf{P}(X)=\prod\limits_{q=0}^{n-1}\left(X-\exp(\tfrac{2\pi i}{n}\left(\alpha_{q}-q-\tfrac{1}{2}\right))\right),$
so that $\textbf{P}\left(\exp(\tfrac{2\pi i}{n}z)\right)=0$ when $z\equiv\alpha_{q}-q-\frac{1}{2}\pmod{n}$. It follows from the definition that the only root of $\textbf{P}\left(\exp(\frac{2\pi i}{n}(z-q-\frac{1}{2}))\right)$ between $-\frac{1}{2}+C$ and $\frac{1}{2}+C$ is at $z=\alpha_{q}$. Moreover, the following lemma about $\textbf{P}(X)$ holds.
Lemma 1.4.1.
@@
For each root $\exp(\frac{2\pi i}{n}(\alpha_{q}-q-\frac{1}{2}))$ of $\textbf{P}(X)$, the complex number
$\exp(-\tfrac{2\pi i}{n}\left(\alpha_{q}-q-\tfrac{1}{2}\right))$ $=\exp(\tfrac{2\pi i}{n}\left(\alpha_{-q-1}+q+\tfrac{1}{2}\right))$
$=\exp(\tfrac{2\pi i}{n}\left(\alpha_{-q-1}-(-q-1)-\tfrac{1}{2}\right))$
@@
$w_{0}(z)$ $=\exp(\tfrac{2\pi i}{n}z)\cdot\frac{G(\exp(\frac{2\pi i}{n}z))}{\textbf{P}(\exp(\frac{2\pi i}{n}z))},$
$w_{q}(z)$ $=w_{0}(z-q),$
where $G$ is a polynomial of degree at most $n-2$ such that $G(X)=X^{n-2}G(X^{-1})$, and let $C$ be a good contour. All traces $T\colon H_{c}\rightarrow\mathbb{C}$ are given by
@@
on $\mathbb{C}[k]e_{q}$, and extended to $H_{c}$ using Corollary 1.3.3 and the values of $T$ on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$.
@@
[…]
@@
    $(z-\alpha_{q})w_{q}(z-\frac{1}{2})$ is holomorphic for $z$ between $-\frac{1}{2}+C$ and $\frac{1}{2}+C$.
Indeed, Property (0) holds because $\textbf{P}(X)$ is not divisible by $X$ and $G$ has degree $\leq n-2$, and Property (0) holds because the only root of $\textbf{P}(\exp(\frac{2\pi i}{n}(z-\frac{1}{2}-q)))$ between $-\frac{1}{2}+C$ and $\frac{1}{2}+C$ is at $z=\alpha_{q}$, and the function $\frac{z-\alpha_{q}}{\exp(\frac{2\pi i}{n}z)-\exp(\frac{2\pi i}{n}\alpha_{q})}$ is holomorphic at $z=\alpha_{q}$.
We now check Conditions (0) and (0) of Theorem 1.3.2. First,
$T\left(\left(k+\tfrac{1}{2}-\alpha_{q}\right)R\left(k+\tfrac{1}{2}\right)e_{q}\right)-T\left(\left(k-\tfrac{1}{2}-\alpha_{q}\right)R\left(k-\tfrac{1}{2}\right)e_{q+1}\right)$
$=\int_{C}\left(z+\tfrac{1}{2}-\alpha_{q}\right)R(z+\tfrac{1}{2})w_{q}(z)\mathrm{d}z-\int_{C}\left(z-\tfrac{1}{2}-\alpha_{q}\right)R\left(z-\tfrac{1}{2}\right)w_{q+1}(z)\mathrm{d}z$
$=\int_{\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\left(z-\tfrac{1}{2}\right)\mathrm{d}z-\int_{-\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q+1}\left(z+\tfrac{1}{2}\right)\mathrm{d}z$
$=\int_{\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\left(z-\tfrac{1}{2}\right)\mathrm{d}z-\int_{-\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\left(z-\tfrac{1}{2}\right)\mathrm{d}z.$
Letting $U$ be the region between $-\frac{1}{2}+C$ and $\frac{1}{2}+C$, this is equal to
@@
which is equal to 0 because $(z-\alpha_{q})R(z)w_{q}(z-\frac{1}{2})$ is holomorphic in $U$ by Property (0) of $w_{q}$ listed at the beginning of the proof. Thus Condition (0) of Theorem 1.3.2 holds.
@@
$w_{0}(z)$ $=\exp(\tfrac{2\pi i}{n}z)\cdot\frac{G\left(\exp(\frac{2\pi i}{n}z)\right)}{\textbf{P}\left(\exp(\frac{2\pi i}{n}z)\right)}$
$=\exp(\tfrac{2\pi i}{n}z)\cdot\frac{\exp((n-2)\cdot\frac{2\pi i}{n}z)G\left(\exp(-\frac{2\pi i}{n}z)\right)}{\exp(n\cdot\frac{2\pi i}{n}z)\textbf{P}\left(\exp(-\frac{2\pi i}{n}z)\right)}$
$=\exp(-\tfrac{2\pi i}{n}z)\cdot\frac{G\left(\exp(-\frac{2\pi i}{n}z)\right)}{\textbf{P}\left(\exp(-\frac{2\pi i}{n}z)\right)}$
$=w_{0}(-z).$
@@
$T\left(R(k)e_{q}\right)$ $=\int_{C}R(z)w_{q}(z)\mathrm{d}z$
$=\int_{C}R(z)w_{0}(z-q)\mathrm{d}z$
$=\int_{C}R(-z)w_{0}(-z-q)\mathrm{d}z$
$=\int_{C}R(-z)w_{0}(z+q)\mathrm{d}z$
$=\int_{C}R(-z)w_{-q}(z)\mathrm{d}z$
$=T\left(R(-k)e_{-q}\right),$
so Condition (0) of Theorem 1.3.2 holds.
Finally, the dimension of the space of possible polynomials $G$ is $\frac{n}{2}$ and distinct polynomials yield distinct traces, so, taking into account the value of the trace on the conjugacy classes $g^{2i}h$ and $g^{2i+1}h$, the dimension of the space of traces described by the theorem is $\frac{n}{2}+2$, which is equal to the dimension of the space of all traces by Proposition 1.3.4. ∎
The following theorem reduces the analytic formula of Theorem 1.4.2 to an integral over the complex line.
Theorem 1.4.3.
Define $w_{q}$ as in Theorem 1.4.2. All traces $T\colon H_{c}\rightarrow\mathbb{C}$ are of the form
$T(R(k)e_{q})=\int_{i\mathbb{R}}R(z)w_{q}(z)\mathrm{d}z+\Phi_{q}(R),$
@@
$\Phi_{q}(R)=\sum_{\begin{subarray}{c}i,j\in\mathbb{Z},\\ i\neq q\end{subarray}}a_{q,i,j}R(\alpha_{i}+q-i+nj-\tfrac{1}{2})$
@@
In the case when all $\alpha_{q}$ satisfy $|\operatorname{Re}(\alpha_{q})|<\frac{1}{2}$, the good contour $C$ may be taken to be $i\mathbb{R}$, proving the theorem. Otherwise, we have
$T(R(k)e_{q})$ $=\int_{C}R(z)w_{q}(z)\mathrm{d}z$
$=\int_{i\mathbb{R}}R(z)w_{q}(z)\mathrm{d}z+\sum_{z_{0}}\operatorname{Res}_{z_{0}}(Rw_{q}),$
where the sum is over all $z_{0}$ between $i\mathbb{R}$ and $C$ where $w_{q}(z)$ has a singularity. In particular, these $z_{0}$ are of the form $\alpha_{i}+q-i+nj-\frac{1}{2}$, with $i\in[0,n-1]$ and $j\in\mathbb{Z}$. The value $\operatorname{Res}_{z_{0}}(Rw_{q})$ is proportional to $R(z_{0})$, proving that the trace $T$ has the form claimed.
If $\Phi_{0}$ is identically zero, then, for all polynomials $R$, the sum $\sum_{z_{0}}\operatorname{Res}_{z_{0}}(Rw_{0})$ over the poles $z_{0}$ of $w_{q}(z)$ between $i\mathbb{R}$ and $C$ is always zero. Taking $R$ to have arbitrary values on the poles (which is possible because the parameters $\alpha_{q}$ are generic, and hence the values $\alpha_{i}+q-i+nj$ are distinct), we find that all residues of these poles are zero. But this implies that there are no poles of $w_{0}$ between $i\mathbb{R}$ and $C$, which is only possible if, for all complex numbers $z_{0}$ of the form $\alpha_{i}-i+nj-\frac{1}{2}$, the monomial $(z-\exp(\frac{2\pi i}{n}z_{0}))$ divides the polynomial $G$ of Theorem 1.4.2. This monomial cancels out the corresponding divisor of the denominator P of $w_{0}$, reducing the trace to one over $P_{\circ}$, as claimed. ∎
Remark 1.4.4.
In the case $q=0$, the poles $\alpha_{i}-i+nj-\frac{1}{2}$ are symmetric about the origin because $\alpha_{i}=-\alpha_{-i-1}$ by Lemma 1.2.3. As was shown in the proof of Theorem 1.4.2, $w_{0}$ is even, so $\operatorname{Res}_{\alpha_{i}-i+nj-\frac{1}{2}}(w_{0})=\operatorname{Res}_{\alpha_{-i-1}+i-nj+\frac{1}{2}}(w_{0})$. It follows that $a_{0,i,j}=a_{0,-i-1,-j}$.
1.5 Positive Traces

### v2

@@ 4 Characterization of traces
$R_{n-1}\bigl(X+n-\tfrac{1}{2}\bigr)-R_{n-1}\bigl(X-\tfrac{1}{2}\bigr)=\sum_{i=0}^{n-1}P_{i}(X+i).$
Then $R_{0}$, …, $R_{n-2}$ are uniquely defined by the first n1n-1 components of the equation, and the resulting solution is valid because the sum $\sum_{i=0}^{n-1}\bigl(R_{i}\bigl(X+i+\frac{1}{2}\bigr)-R_{i-1}\bigl(X+i-\frac{1}{2}\bigr)\bigr)$ is equal to $R_{n-1}\bigl(X+n-\frac{1}{2}\bigr)-R_{n-1}\bigl(X-\frac{1}{2}\bigr)$}, which equals $\sum_{i=0}^{n-1}P_{i}(X+i)$, as needed. We also note that the kernel of $\varphi$ is $\{(z,\dots,z)\mid z\in\mathbb{C}\}$.
Letting $U=\{(R_{0}(X),\dots,R_{n-1}(X))\mid R_{i}(X)=R_{-i-1}(-X)\}$, we claim that the preimage of $W$ under $\varphi$ is $U$. Indeed, $\varphi(U)\subseteq W$ because for any element $(R_{0}(X),\dots,R_{n-1}(X))\in U$, the $i$th component of $\varphi(R_{0}(X),\dots,R_{n-1}(X))$ is
$R_{i}\bigl(X+\tfrac{1}{2}\bigr)-R_{i-1}\bigl(X-\tfrac{1}{2}\bigr)=R_{-i-1}\bigl(-X-\tfrac{1}{2}\bigr)-R_{-i}\bigl(-X+\tfrac{1}{2}\bigr),$
which is equal to the negative of the (ni)(n-i)th component. Conversely, for any element $(R_{0}(X),\allowbreak\dots,R_{n-1}(X))$ mapped to $W$ by $\varphi$, the element of $V$ with $i$th component equal to
$\frac{R_{i}(X)+R_{-i-1}(-X)}{2}$
belongs to $U$ and is also mapped to $W$ by $\varphi$. Therefore, $V/\langle\varphi(I),W\rangle\cong V/\langle I,U\rangle$.
Let $\psi\colon V\to V/I\cong\mathbb{C}^{n}$ be the quotient map given by mapping $(R_{0}(X),\dots,R_{n-1}(X))$ to $(R_{0}(\alpha_{0}),\dots,R_{n-1}(\alpha_{n-1}))$. Because $\alpha_{i}=-\alpha_{-i-1}$, $\psi(U)$ is the set $S$ of vectors $(v_{0},\dots,v_{n-1})$ such that $v_{i}=v_{-i-1}$ for all $i$. Therefore, $V/\langle I,U\rangle\cong\mathbb{C}^{n}/S$, the dimension of which is $\frac{n}{2}$, as desired. ∎
5 Analytic formula for traces
We now give an analytic description of traces using the notion of a good contour from [18]. In particular, similar to [18, Definition 3.6], we say that a non-self-intersecting curve $C$ with exactly one point in the set $\mathbb{R}+ri$ for all real $r$ is a good contour if the following three conditions hold:
  1. $(1)$
    There exist real $r>0$ and $a$ such that, letting $B_{r}$ be the disk of radius $r$ centered at the origin, $C\setminus B_{r}$ coincides with $(a+{\rm i}\mathbb{R}_{>0})\cup(-a+{\rm i}\mathbb{R}_{<0})\setminus B_{r}$. This property ensures that the notions of “to the left of CC” and “to the right of C” are well-defined.
  2. $(2)$
@@
  3. $(3)$
    The path $C$ is symmetric about the origin, that is, for every complex number $z$ on $C$, z-z is on $C$.
Because the parameter $c$ is assumed to be generic, no two $\alpha_{q}$ have the same imaginary part, so a curve satisfying condition (2) exists. Also, since $\alpha_{q}=-\alpha_{-q-1}$ by Lemma 3.3, the curve may be taken to also satisfy condition (3). Finally, since there are finitely many $\alpha_{q}$, the curve may also be taken to satisfy condition (1), so a good contour is guaranteed to exist.
@@ 3. $(3)$
$\textbf{P}(X)=\prod\limits_{q=0}^{n-1}\bigl(X-\exp\bigl(\tfrac{2\pi{\rm i}}{n}\bigl(\alpha_{q}-q-\tfrac{1}{2}\bigr)\bigr)\bigr),$
so that $\textbf{P}\bigl(\exp\bigl(\tfrac{2\pi{\rm i}}{n}z\bigr)\bigr)=0$ when $z\equiv\alpha_{q}-q-\frac{1}{2}\pmod{n}$. It follows from the definition that the only root of $\textbf{P}\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}\bigl(z-q-\frac{1}{2}\bigr)\bigr)\bigr)$ between $\frac{1}{2}+C$ and $\frac{1}{2}+C$ is at $z=\alpha_{q}$. Moreover, the following lemma about $\textbf{P}(X)$ holds.
Lemma 5.1.
@@
For each root $\exp\bigl(\frac{2\pi{\rm i}}{n}\bigl(\alpha_{q}-q-\frac{1}{2}\bigr)\bigr)$ of $\textbf{P}(X)$, the complex number
$\exp\bigl(-\tfrac{2\pi{\rm i}}{n}\bigl(\alpha_{q}-q-\tfrac{1}{2}\bigr)\bigr)={}$ $\exp\bigl(\tfrac{2\pi{\rm i}}{n}\bigl(\alpha_{-q-1}+q+\tfrac{1}{2}\bigr)\bigr)=\exp\bigl(\tfrac{2\pi{\rm i}}{n}\bigl(\alpha_{-q-1}-(-q-1)-\tfrac{1}{2}\bigr)\bigr)$
@@
$w_{0}(z)=\exp\bigl(\tfrac{2\pi{\rm i}}{n}z\bigr)\cdot\frac{G\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}{\textbf{P}\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}z\bigr)\bigr)},\qquad w_{q}(z)=w_{0}(z-q),$
where $G$ is a polynomial of degree at most n2n-2 such that $G(X)=X^{n-2}G\bigl(X^{-1}\bigr)$, and let $C$ be a good contour. All traces $T\colon H_{c}\rightarrow\mathbb{C}$ are given by
@@
on $\mathbb{C}[k]e_{q}$, and extended to $H_{c}$ using Corollary 4.3 and the values of $T$ on the conjugacy classes $\big\{g^{2i}h\mid i\in\mathbb{Z}\big\}$} and $\big\{g^{2i+1}h\mid i\in\mathbb{Z}\big\}$.
@@
    $w_{q}(z)$ decays exponentially when $\operatorname{Im}(z)$ approaches $\pm\infty$;
@@
    $(z-\alpha_{q})w_{q}\bigl(z-\frac{1}{2}\bigr)$ is holomorphic for $z$ between $\frac{1}{2}+C$ and $\frac{1}{2}+C$.
Indeed, Property (0) holds because $\textbf{P}(X)$ is not divisible by $X$ and $G$ has degree $\leq n-2$, and Property (0) holds because the only root of $\textbf{P}\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}\bigl(z-\frac{1}{2}-q\bigr)\bigr)\bigr)$ between $\frac{1}{2}+C$ and $\frac{1}{2}+C$ is at $z=\alpha_{q}$, and the function
$\frac{z-\alpha_{q}}{\exp\bigl(\frac{2\pi{\rm i}}{n}z\bigr)-\exp\bigl(\frac{2\pi{\rm i}}{n}\alpha_{q}\bigr)}$
is holomorphic at $z=\alpha_{q}$.
We now check conditions (0) and (0) of Theorem 4.2. First,
$T\bigl(\bigl(k+\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k+\tfrac{1}{2}\bigr)e_{q}\bigr)-T\bigl(\bigl(k-\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(k-\tfrac{1}{2}\bigr)e_{q+1}\bigr)$
$\qquad=\int_{C}\bigl(z+\tfrac{1}{2}-\alpha_{q}\bigr)R(z+\tfrac{1}{2})w_{q}(z)\mathrm{d}z-\int_{C}\bigl(z-\tfrac{1}{2}-\alpha_{q}\bigr)R\bigl(z-\tfrac{1}{2}\bigr)w_{q+1}(z)\mathrm{d}z$
$\qquad=\int_{\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\bigl(z-\tfrac{1}{2}\bigr)\mathrm{d}z-\int_{-\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q+1}\bigl(z+\tfrac{1}{2}\bigr)\mathrm{d}z$
$\qquad=\int_{\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\bigl(z-\tfrac{1}{2}\bigr)\mathrm{d}z-\int_{-\frac{1}{2}+C}(z-\alpha_{q})R(z)w_{q}\bigl(z-\tfrac{1}{2}\bigr)\mathrm{d}z.$
Letting $U$ be the region between $\frac{1}{2}+C$ and $\frac{1}{2}+C$, this is equal to
@@
which is equal to 0 because $(z-\alpha_{q})R(z)w_{q}\bigl(z-\frac{1}{2}\bigr)$ is holomorphic in $U$ by Property (0) of $w_{q}$ listed at the beginning of the proof. Thus condition (0) of Theorem 4.2 holds.
@@
$w_{0}(z)={}$ $\exp\bigl(\tfrac{2\pi{\rm i}}{n}z\bigr)\cdot\frac{G\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}{\textbf{P}\bigl(\exp\bigl(\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}=\exp\bigl(\tfrac{2\pi{\rm i}}{n}z\bigr)\cdot\frac{\exp\bigl((n-2)\cdot\frac{2\pi{\rm i}}{n}z\bigr)G\bigl(\exp\bigl(-\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}{\exp\bigl(n\cdot\frac{2\pi{\rm i}}{n}z\bigr)\textbf{P}\bigl(\exp\bigl(-\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}$
$={}$ $\exp\bigl(-\tfrac{2\pi{\rm i}}{n}z\bigr)\cdot\frac{G\bigl(\exp\bigl(-\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}{\textbf{P}\bigl(\exp\bigl(-\frac{2\pi{\rm i}}{n}z\bigr)\bigr)}=w_{0}(-z).$
@@
$T(R(k)e_{q})={}$ $\int_{C}R(z)w_{q}(z)\mathrm{d}z=\int_{C}R(z)w_{0}(z-q)\mathrm{d}z=\int_{C}R(-z)w_{0}(-z-q)\mathrm{d}z$
$={}$ $\int_{C}R(-z)w_{0}(z+q)\mathrm{d}z=\int_{C}R(-z)w_{-q}(z)\mathrm{d}z=T(R(-k)e_{-q}),$
so condition (0) of Theorem 4.2 holds.
Finally, the dimension of the space of possible polynomials $G$ is $\frac{n}{2}$ and distinct polynomials yield distinct traces, so, taking into account the value of the trace on the conjugacy classes $\big\{g^{2i}h\mid i\in\mathbb{Z}\big\}$ and $\big\{g^{2i+1}h\mid i\in\mathbb{Z}\big\}$, the dimension of the space of traces described by the theorem is $\frac{n}{2}+2$, which is equal to the dimension of the space of all traces by Proposition 4.4. ∎
The following theorem reduces the analytic formula of Theorem 5.2 to an integral over the complex line.
Theorem 5.3.
Define $w_{q}$ as in Theorem 5.2. All traces $T\colon H_{c}\rightarrow\mathbb{C}$ are of the form
$T(R(k)e_{q})=\int_{{\rm i}\mathbb{R}}R(z)w_{q}(z)\mathrm{d}z+\Phi_{q}(R),$
@@
$\Phi_{q}(R)=\sum_{\begin{subarray}{c}i,j\in\mathbb{Z},\\ i\neq q\end{subarray}}a_{q,i,j}R\bigl(\alpha_{i}+q-i+nj-\tfrac{1}{2}\bigr)$
@@
In the case when all $\alpha_{q}$ satisfy $|{\operatorname{Re}(\alpha_{q})}|<\frac{1}{2}$, the good contour $C$ may be taken to be $\rm i$$\mathbb{R}$, proving the theorem. Otherwise, we have
$T(R(k)e_{q})={}$ $\int_{C}R(z)w_{q}(z)\mathrm{d}z=\int_{{\rm i}\mathbb{R}}R(z)w_{q}(z)\mathrm{d}z+\sum_{z_{0}}\operatorname{Res}_{z_{0}}(Rw_{q}),$
where the sum is over all $z_{0}$ between $\rm i$$\mathbb{R}$ and $C$ where $w_{q}(z)$ has a singularity. In particular, these $z_{0}$ are of the form $\alpha_{i}+q-i+nj-\frac{1}{2}$, with $i\in[0,n-1]$ and $j\in\mathbb{Z}$. The value $\operatorname{Res}_{z_{0}}(Rw_{q})$ is proportional to $R(z_{0})$, proving that the trace $T$ has the form claimed.
If $\Phi_{0}$ is identically zero, then, for all polynomials $R$, the sum $\sum_{z_{0}}\operatorname{Res}_{z_{0}}(Rw_{0})$ over the poles $z_{0}$ of $w_{q}(z)$ between $\rm i$$\mathbb{R}$ and $C$ is always zero. Taking $R$ to have arbitrary values on the poles (which is possible because the parameters $\alpha_{q}$ are generic, and hence the values $\alpha_{i}+q-i+nj$ are distinct), we find that all residues of these poles are zero. But this implies that there are no poles of $w_{0}$ between $\rm i$$\mathbb{R}$ and $C$, which is only possible if, for all complex numbers $z_{0}$ of the form $\alpha_{i}-i+nj-\frac{1}{2}$}, the monomial $\bigl(z-\exp\bigl(\frac{2\pi{\rm i}}{n}z_{0}\bigr)\bigr)$ divides the polynomial $G$ of Theorem 5.2. This monomial cancels out the corresponding divisor of the denominator P of $w_{0}$, reducing the trace to one over $P_{\circ}$, as claimed. ∎
Remark 5.4.
In the case $q=0$, the poles $\alpha_{i}-i+nj-\frac{1}{2}$ are symmetric about the origin because $\alpha_{i}=-\alpha_{-i-1}$ by Lemma 3.3. As was shown in the proof of Theorem 5.2, $w_{0}$ is even, so $\operatorname{Res}_{\alpha_{i}-i+nj-\frac{1}{2}}(w_{0})=\operatorname{Res}_{\alpha_{-i-1}+i-nj+\frac{1}{2}}(w_{0})$. It follows that $a_{0,i,j}=a_{0,-i-1,-j}$.
6 Positive traces